# A Robust Mixed-Effects Bandit Algorithm for Assessing Mobile Health Interventions

### v3

@@ F.2 Preliminaries
In the limit as $k\rightarrow\infty$
$\tilde{\sigma}^{2}_{i,t}\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $\leq\pi_{i,t}(0|s)\textrm{Var}(\epsilon_{i,t})+(1-\pi_{i,t}(0|s))\textrm{Var}(\epsilon_{i,t})$
$=\textrm{Var}(\epsilon_{i,t})$
In the next remark, we show what the variance would be if we did not use DML and estimate $f_{i,t}\approx r_{i,t}$, but used only the inverse propensity weighted observed reward as the pseudo-reward. This was done in Greenewald et al. (2017). In this case, there are terms dependent on the mean reward that do not vanish as the number of stages goes to infinity.
Remark 2.
If we instead used as our pseudo-reward the inverse propensity weighted observed reward
$\tilde{R}_{i,t}(s,\bar{a})$ $=\frac{R_{it}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}$
this would be unbiased with variance
$v_{k}^{2}\equiv\frac{r_{i,t}(s,\bar{a})^{2}+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{r_{i,t}(s,0)^{2}+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}-\Delta_{i,t}(s,\bar{a})^{2}$
@@
The unbiasedness is clear from our proof of Lemma 2. For the variance,
$\textrm{Var}\left(\frac{R_{it}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\right)$ $=\mathbb{E}\left[\frac{(r_{i,t}(s,\bar{a})+\epsilon_{i,t})^{2}}{1-\pi_{i,t}(0|s)}+\frac{(r_{i,t}(s,0)+\epsilon_{i,t})^{2}}{\pi_{i,t}(0|s)}\right]-\Delta_{i,t}(s,\bar{a})^{2}$
$=\frac{r_{i,t}(s,\bar{a})^{2}+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{r_{i,t}(s,0)^{2}+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}-\Delta_{i,t}(s,\bar{a})^{2}$
$=\frac{r_{i,t}(s,\bar{a})^{2}+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{r_{i,t}(s,0)^{2}+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad-(r_{i,t}(s,\bar{a})-r_{i,t}(s,0))^{2}$
$=\frac{r_{i,t}(s,\bar{a})^{2}+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}-r_{i,t}(s,\bar{a})^{2}+\frac{r_{i,t}(s,0)^{2}+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad-r_{i,t}(s,0)^{2}+2r_{i,t}(s,\bar{a})r_{i,t}(s,0)$
$=\frac{r_{i,t}(s,\bar{a})^{2}\pi_{i,t}(0,s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{r_{i,t}(s,0)^{2}[1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad+2r_{i,t}(s,\bar{a})r_{i,t}(s,0)$
Here we collect some important results. We first adapt an important concentration inequality for regularized least-squares estimates. Our proof follows the same basic strategy as Abbasi-Yadkori et al. (2011), but with modifications due to (a) the use of weighted least squares (b) the use of pseudo-rewards to estimate differential rewards (c) replacing the scaled diagonal regularization with Laplacian regularization.
Lemma 6 (Adapted from Theorem 2 in Abbasi-Yadkori et al. (2011)).
Let $\hat{\theta}_{k}$ be the stage $k$ regularized least squares (RLS) estimate from Algorithm 1 and $\theta^{*}$ the ground truth. For any $\delta>0$, w.p. at least $1-\delta$ the estimates $\{\theta_{k}\}_{k=0}^{\infty}$ satisfies for any $\{x_{k}\}_{k=0}^{\infty}$,
$|x_{k}^{\top}(\hat{\theta}_{k}-\theta^{*})|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right),$ (16)
where $v_{k}^{2}$ is the variance factor for the difference between the pseudo-reward and its mean at stage $k$. In particular, setting $x_{k}=V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})$ implies
$\|\hat{\theta}_{k}-\theta^{*}\|_{V_{k-1}}$ $\leq v_{k}\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB$
holds w.p. at least $1-\delta$ for all $k\geq 1$.
@@
Let $m_{i,t}=\tilde{\sigma}_{i,t}\phi(\textbf{x}_{i,t})$ and $\rho_{i,t}=\tilde{\sigma}_{i,t}[\tilde{R}_{i,t}^{f}(s,\bar{a})-\mathbb{E}[\tilde{R}_{i,t}^{f}|s,\bar{a}]]$. Further let
$\xi_{k}$ $\equiv\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}^{2}[\tilde{R}_{i,t}^{f}(s,\bar{a})-\mathbb{E}[\tilde{R}_{i,t}^{f}|s,\bar{a}]]\phi(\textbf{x}_{i,t})$
$=\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}[\tilde{R}_{i,t}^{f}(s,\bar{a})-\mathbb{E}[\tilde{R}_{i,t}^{f}|s,\bar{a}]]m_{i,t}$
$=\sum_{(i,t)\in\mathcal{O}_{k}}m_{i,t}\rho_{i,t}$
Then noting that $V_{k}=\sum_{(i,t)\in\mathcal{O}_{k}}m_{i,t}m_{i,t}^{\top}+V_{0}$ and letting $W_{k}$ be the diagonal matrix of weights $\tilde{\sigma}_{i,t}^{2}$, we have
$\hat{\theta}_{k}$ $=V_{k-1}^{-1}b_{k}$
$=V_{k-1}^{-1}(\xi_{k}+\Phi_{k}^{\top}W_{k}\mathbb{E}[R_{k}^{f_{k}}|\bar{a},s])$
@@
$=V_{k-1}^{-1}\xi_{k}+V_{k-1}^{-1}\Phi_{k}^{\top}W_{k}\Phi_{k}\theta^{*}$
$=V_{k-1}^{-1}\xi_{k}+V_{k-1}^{-1}(\Phi_{k}^{\top}W_{k}\Phi_{k}+V_{0})\theta^{*}-V_{k-1}^{-1}V_{0}\theta^{*}$
$=V_{k-1}^{-1}\xi_{k}+\theta^{*}-V_{k-1}^{-1}V_{0}\theta_{k}^{*}$
and thus
$\hat{\theta}_{k}-\theta^{*}$ $=V_{k-1}^{-1}(\xi_{k}-V_{0}\theta^{*})$
which gives
$|x_{k}^{\top}\hat{\theta}_{k}-x_{k}^{\top}\theta^{*}|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}(\|\xi_{k}\|_{V_{k-1}^{-1}}+\|V_{0}\theta*\|_{V_{k-1}^{-1}})$
Now since $\xi_{k}$ is sub-Gaussian with variance factor $v_{k}^{2}$, by Theorem 1 in Abbasi-Yadkori et al. (2011), w.p. $1-\delta$,
$\|\xi_{k}\|_{V_{k-1}^{-1}}^{2}$ $\leq 2v_{k}^{2}\log\left(\frac{\textrm{det}(V_{k-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)$
Further note
$\|V_{0}\theta^{*}\|_{V_{k-1}^{-1}}^{2}$ $=\theta_{k}^{*\top}V_{0}^{\top}V_{k-1}^{-1}V_{0}\theta^{*}$
$\leq\|V_{0}^{\top}V_{k-1}^{-1}V_{0}\|_{2}\|\theta^{*}\|_{2}^{2}$
$\leq\|V_{0}\|^{2}_{2}\|V_{k-1}^{-1}\|_{2}\|\theta^{*}\|_{2}^{2}$
$\leq\lambda_{\textrm{max}}(V_{0})\|\theta^{*}\|_{2}^{2}$
$\leq\lambda_{\textrm{max}}(V_{0})k^{2}B^{2}$
and thus
$|x_{k}^{\top}\hat{\theta}_{k}-x_{k}^{\top}\theta^{*}|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right).$
Finally, setting $x_{k}=V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})$ implies
$\|\hat{\theta}_{k-1}-\theta^{*}\|^{2}_{V_{k-1}}$ $=(\hat{\theta}_{k-1}-\theta^{*})^{\top}V_{k-1}^{\top}(\hat{\theta}_{k-1}-\theta^{*})$
$=|x_{k}^{\top}\hat{\theta}_{k}-x_{k}^{\top}\theta^{*}|$
$\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right)$
$=\|V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right)$
$=\|\hat{\theta}_{k-1}-\theta^{*}\|_{V_{k-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right)$
and dividing both sides by $\|\hat{\theta}_{k-1}-\theta^{*}\|_{V_{k-1}}$ implies
$\|\hat{\theta}_{k}-\theta^{*}\|_{V_{k-1}}$ $\leq v_{k}\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB$
Corollary 5.
If Assumption 3, then for any $\delta>0$, there exists $C>0$ s.t. w.p. at least $1-\delta$ the estimates $\{\hat{\theta}_{k}\}_{k=0}^{\infty}$ in Algorithm 2 satisfies for any $\{x_{k}\}_{k=0}^{\infty}$,
$|x_{k}^{\top}(\hat{\theta}_{k}-\theta^{*})|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(\left(\frac{C}{k^{1/2}}+\sigma^{2}c^{2}\right)\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta/2}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right),$ (17)
In particular, setting $x_{k}=V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})$ implies
$\|\hat{\theta}_{k}-\theta^{*}\|_{V_{k-1}}$ $\leq\left(\frac{C}{k^{1/2}}+\sigma^{2}c^{2}\right)\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta/2}\right)}+\lambda_{\max}(V_{0})^{1/2}kB$
holds w.p. at least $1-\delta$ for all $k\geq 1$.
@@
Use Corollary 3 and Lemma 6, each with $\delta/2$. Then w.p. at least $1-\delta$ the result holds. ∎
We next state a slightly modified form of a standard result of RLS (Lemma 11 in Abbasi-Yadkori et al. (2011)) that helps to guarantee that the prediction error is cumulatively small. This bounds the sum of quadratic forms where the matrix is the inverse Gram matrix and the arguments are the feature vectors. We use such terms to construct a martingale in the regret bound so that we can bound such terms and the martingale.
Proposition 1.
Let $\lambda\geq 1$ and $\gamma\geq 1$. For any arbitrary sequence $(x_{i,t})_{(i,t)\in\mathcal{O}_{k}}$, let
$V_{k+1}\equiv\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}^{2}\phi(x_{i,t})\phi(x_{i,t})^{\top}+V_{0},$
be the regularized Gram matrix. Then
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}\leq 2c\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right).$
where $c$ is a constant such that $0<\frac{1}{c}<\tilde{\sigma}_{i,t}^{2}\forall i,t\in\mathbb{N}$. Further,
$\log\det(V_{K+1})$ $\leq 2Kd\log\left(\left[\gamma+\lambda M+\frac{K+1}{8d}\right]\right)$

### v4

@@ E.3 Derivation of Confidence Sets
We can apply Lemma 7 and Assumption 8 to bound $\|C_{i,t}V_{i,t}^{-1}V_{0,k}\theta_{k}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}$. Assumption 8 implies that there exists a function $\zeta(\delta):(0,1)\to\mathbb{R}$ such that
$\sup_{(i,t)\in\mathcal{O}_{K}}\|C_{i,t}V_{i,t}^{-1}V_{0,k}\theta_{k}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}\leq\sup_{(i,t)\in\mathcal{O}_{K}}\|V_{0,k}\theta_{k}^{\star}\|_{V_{i,t}^{-1}}\leq\zeta(\delta)\max\{\log^{3/4}(K),1\}$
with probability at least $1-\delta$. This function corresponds with the hyperparameter $\zeta$ given in the main paper. The purpose of the $\max$ operation is to avoid having a bound of zero for $K=1$.
We confirmed in simulation that Assumption 8 assumption is reasonable. The left panel of Figure 13 provides empirical evidence that $\|V_{0,k}\theta_{k}^{\star}\|_{V_{i,t}^{-1}}/\log^{3/4}(k)$ is bounded in probability in a simulation with values of $k$ ranging from 1 to 200. We ran 400 repetitions with an intercept-only model. Similarly, the right panel suggests that $\sup_{(i,t)\in\mathcal{O}_{K}}\|V_{0,k}\theta_{k}^{\star}\|_{V_{i,t}^{-1}}/\log^{3/4}(K)$ is bounded as assumed. We conjecture that the following assumption is a sufficient condition (in combination with the other assumptions listed in the main paper) for Assumption 8 to hold.
Figure 13: (left) Simulated values of $\|V_{0,k}\theta_{k}^{\star}\|_{V_{i,t}^{-1}}/\log^{3/4}(k)$. (right) Simulated values of $\sup_{(i,t)\in\mathcal{O}_{K}}\|V_{0,k}\theta_{k}^{\star}\|_{V_{i,t}^{-1}}/\log^{3/4}(K)$. The plots show evidence that these terms are bounded as required by Assumption 8.
Assumption 11.
The random effects $\theta_{i}^{\text{user}}$, $\theta_{t}^{\text{time}}$ for $i,t=1,2,\ldots$ are distributed such that
$\frac{1}{\sqrt{K}}\sum_{k=1}^{K}\theta_{k}^{\text{user}}O_{p}(1),\quad\frac{1}{\sqrt{K}}\sum_{k=1}^{K}\theta_{k}^{\text{time}}=O_{p}(1).$
So, in other words, it would be sufficient for the normalized sums above to converge in distribution. This effectively requires these normalized sums to be approximately mean zero. This requirement is reasonable given the presence of $\theta^{\text{shared}}$ in the model.
We now bound the first term from Lemma 6.
Lemma 8.
Assuming $\gamma\geq 1$, we have:
$\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})\leq\left(\frac{3}{\gamma}\right)^{d}$
@@
By definition, we have:
[…]
Taking the determinant of both sides:
$\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})=\det(C_{i,t}V_{i,t}^{-1}C_{i,t}^{{\top}})$
Using the property that the determinant of a matrix is less than or equal to the product of its eigenvalues, and letting $\lambda_{j}(\cdot)$ denote the $j$-th largest eigenvalue, we have:
$\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})$ $\leq\prod_{j=1}^{d}\lambda_{j}(C_{i,t}V_{i,t}^{-1}C_{i,t}^{{\top}})$
$\leq\lambda_{\text{max}}(C_{i,t}V_{i,t}^{-1}C_{i,t}^{{\top}})^{d}$
Now letting $\sigma_{1}$ be the largest singular value of a matrix,
$\lambda_{\text{max}}(C_{i,t}V_{i,t}^{-1}C_{i,t}^{{\top}})\leq\sigma_{1}(C_{i,t})^{2}\lambda_{\text{max}}(V_{i,t}^{-1})$
Since $C_{i,t}=(1_{3}^{{\top}}\otimes I_{d})\underline{\hbox{\set@color$C$}}_{i,t}$, we have that
$\sigma_{1}(C_{i,t})$ $=\sigma_{1}((1_{3}^{{\top}}\otimes I_{d})\underline{\hbox{\set@color$C$}}_{i,t})$
$\leq\sigma_{1}(1_{3}^{{\top}}\otimes I_{d})\sigma_{1}(\underline{\hbox{\set@color$C$}}_{i,t})$
$=\sigma_{1}(1_{3}^{\top})\sigma_{1}(I_{d})\sigma_{1}(\underline{\hbox{\set@color$C$}}_{i,t})$
$=\sqrt{3}\cdot 1\cdot 1$
$=\sqrt{3}$
so that $\sigma_{1}(C_{i,t})^{2}=3$. Therefore,
$\lambda_{\text{max}}(C_{i,t}V_{i,t}^{-1}C_{i,t}^{{\top}})$ $\leq 3\lambda_{\text{max}}(V_{i,t}^{-1})$
$=\frac{3}{\lambda_{\min}(V_{i,t})}$
$\leq\frac{3}{\gamma}$
and thus
$\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})$ $\leq\left(\frac{3}{\gamma}\right)^{d}.$
Lemma 9.
$\text{det}(\Lambda_{i,t}^{0})$ $\geq 3^{d}\left(\frac{\lambda_{\min}(V_{0,k})}{\lambda_{\max}(V_{i,t})^{2}}\right)^{d}$
@@
$\text{det}(\Lambda_{i,t}^{0})$ $\geq\prod_{j=1}^{d}\sigma_{\min}(1_{3}^{\top}\otimes I_{d})^{2}\lambda_{\min}(C_{i,t}V_{i,t}^{-1}V_{0,k}V_{i,t}^{-1}C_{i,t}^{\top})$
$\geq 3^{d}\lambda_{\min}(V_{i,t}^{-1}V_{0,k}V_{i,t}^{-1})^{d}$
$\geq 3^{d}\left(\frac{\lambda_{\min}(V_{0,k})}{\lambda_{\max}(V_{i,t})^{2}}\right)^{d}$
Lemma 10.
Assuming $\gamma\geq 1$,
$\log\frac{\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})}{\det(\Lambda^{0}_{i,t})}$ $\leq 2d\log\left\{\frac{3k(k+1)}{8}+\gamma k+2\lambda e_{k}\right\}.$
Proof.
$\log\frac{\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})}{\det(\Lambda^{0}_{i,t})}$ $\leq d\log\frac{3}{\gamma}-d\log 3+2d\log\lambda_{\max}(V_{i,t})-d\log\lambda_{\min}(V_{0,k})$
$\leq 2d(\log\lambda_{\max}(V_{i,t})-\log\gamma)$
Now we need to upper bound $\lambda_{\max}(V_{i,t})$. Note that
$\lambda_{\max}(V_{i,t})$ $\leq\textrm{tr}\left(V_{i,t}\right)$
$\leq\text{tr}\left(V_{0,k}+\sum_{(i,t)\in\mathcal{O}_{K}}\tilde{\sigma}_{i,t}^{2}\phi(x_{i,t})\phi(x_{i,t})^{\top}\right)$
$\leq\sum_{(i,t)\in\mathcal{O}_{K}}\tilde{\sigma}_{i,t}^{2}\|\phi(x_{i,t})\|_{2}^{2}+\gamma k+2\lambda e_{k}$
$\leq\frac{3}{4}\frac{k(k+1)}{2}+\gamma k+2\lambda e_{k}.$
This then gives us
$\log\frac{\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})}{\det(\Lambda^{0}_{i,t})}$ $\leq 2d\left(\log\frac{3k(k+1)}{8}+\gamma k+2\lambda e_{k}-\log\gamma\right)$
$\leq 2d\log\left\{\frac{3k(k+1)}{8}+\gamma k+2\lambda e_{k}\right\}$
Note that $e_{k}$ produces no more than a logarithmic contribution to the above bound due to Assumption 7. We now define a term, $\beta_{i,t}(\delta_{1},\delta_{2})$, that corresponds with the term $\beta_{i,t}(\delta)$ from the main paper, setting $\delta=2\delta_{1}$ with sufficiently large $\zeta$. However, here we make explicit that the term depends on two probabilities, $\delta_{1}$ and $\delta_{2}$.
Definition 2.
Let $\beta_{i,t}(\delta_{1},\delta_{2})$ be defined as follows:
$\beta_{i,t}(\delta_{1},\delta_{2})\coloneq v\sqrt{2\log\left(\frac{1}{\delta_{1}}\right)+\log\left\{\frac{\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})}{\det(\Lambda_{i,t}^{0})}\right\}}+\zeta(\delta_{2})\max\{\log^{3/4}(K),1\}.$
Note that by Lemma 6 and Assumption 8 we have $\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}\leq\beta_{i,t}(\delta_{1},\delta_{2})$ with probability at least $1-\delta_{1}-\delta_{2}$.
Lemma 11.
Let $\delta_{1}^{\prime}=\delta_{1}/\{K(K+1)\}$. Then
$\sup_{(i,t)\in\mathcal{O}_{K}}\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}\leq\sup_{(i,t)\in\mathcal{O}_{K}}\beta_{i,t}(\delta_{1}^{\prime},\delta_{2})$
with probability at least $1-\delta_{1}/2-\delta_{2}$. Further, $\sup_{(i,t)\in\mathcal{O}_{K}}\beta_{i,t}(\delta_{1}^{\prime},\delta_{2})\leq\beta_{K}(\delta_{1},\delta_{2})$, where $\beta_{K}(\delta_{1},\delta_{2})$ is defined as
$v\sqrt{2\log\left\{\frac{K(K+1)}{\delta_{1}}\right\}+2d\log\left\{\frac{3K(K+1)}{8}+\gamma K+2\lambda e_{K}\right\}}+\zeta(\delta_{2})\max\{\log^{3/4}(K),1\}.$
Proof.
The first statement follows from applying Lemma 6 with a union bound. Note that the second term in Lemma 6 can be bounded uniformly using Lemma 7 and Assumption 8 as follows:
$\|C_{i,t}V_{i,t}^{-1}V_{0,k}\theta_{k}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}\leq\|V_{0,k}\theta_{k}^{\star}\|_{V_{i,t}^{-1}}\leq\zeta(\delta_{2})\max\{\log^{3/4}(K),1\}$
with probability at least $1-\delta_{2}$. The second statement follows from bounding the first term in $\beta_{i,t}(\delta_{1}^{\prime},\delta_{2})$ using Lemma 10:
$v\sqrt{2\log\left\{\frac{K(K+1)}{\delta_{1}}\right\}+\log\frac{\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})}{\det(\Lambda_{i,t}^{0})}}$
@@
Lastly, we define several events and bound the probabilities with which they occur.
Proposition 1.
Let
  • $\delta_{1}^{\prime}=\delta_{1}/\{K(K+1)\}$
  • $\hat{E}_{K}=\{\forall(i,t)\in\mathcal{O}_{K},\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}\leq\beta_{i,t}(\delta_{1}^{\prime},\delta_{2})\}$
  • $\gamma_{i,t}(\delta_{1},\delta_{2})=\beta_{i,t}(\delta_{1},\delta_{2})\sqrt{cd\log(c^{\prime}d/\delta_{1})}$
  • $\tilde{E}_{K}=\{\forall(i,t)\in\mathcal{O}_{K},\|\tilde{\theta}_{i,t}-\check{\theta}_{i,t}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}\leq\gamma_{i,t}(\delta_{1}^{\prime},\delta_{2})\}$
  • $E_{K}=\hat{E}_{K}\cap\tilde{E}_{K}$
Then $\Pr(E_{K})\geq 1-\delta_{1}-\delta_{2}$.
@@
First, Lemma 11 shows that $\Pr(\hat{E}_{k})\geq 1-\delta_{1}/2-\delta_{2}$.
Next, applying the TS sampling distribution and $\tilde{\theta}_{i,t}=\check{\theta}_{i,t}+\beta_{k}(\delta_{1}^{\prime},\delta_{2})\underline{\hbox{\set@color$V$}}_{i,t}^{-1/2}\eta_{i,t}$ where $\eta_{i,t}$ is drawn i.i.d. from $\mathcal{D}^{TS}$ we have
$\Pr\left(\|\tilde{\theta}_{i,t}-\check{\theta}_{i,t}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}\leq\beta_{i,t}(\delta_{1}^{\prime},\delta_{2})\sqrt{cd\log\left(\frac{c^{\prime}d}{\delta_{1}^{\prime}}\right)}\right)$
=\ = $\Pr\left(\|\eta_{i,t}\|\leq\sqrt{cd\log\left(\frac{c^{\prime}d}{\delta_{1}^{\prime}}\right)}\right)$
[…]
by the concentration property in Definition 1. A union bound then gives $\Pr(\tilde{E}_{K})\geq 1-\delta_{1}/2$. Applying a union bound to $\tilde{E}_{K}$ and $\hat{E}_{K}$ yields the conclusion. ∎
E.4 Treatment Probability Estimation Errors
The proof relies on bounding terms of the form $\{\pi_{i,t}(\bar{A}_{i,t}|S_{i,t})-\pi_{i,t}^{\star}(\bar{A}_{i,t}^{\star}|S_{i,t})\}x(S_{i,t},\bar{A}_{i,t})^{\top}\theta_{i,t}^{\star}$. To simplify the proof, we denote these terms as $(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}$. We begin by defining two filtrations.
Definition 2.
We define the filtration $\mathcal{F}_{i,t}^{S}$ as the information accumulated up to decision point $(i,t)$ including the sampled context, that is, $\mathcal{F}_{i,t}^{S}=(\mathcal{F}_{1},\sigma(S_{1,1},\bar{A}_{1,1},A_{1,1},R_{1,1}S_{2,1},\ldots,S_{i,t}))$. Similarly, we define the filtration $\mathcal{F}_{i,t}$ as the information accumulated up to decision point $(i,t)$ before sampling $S_{i,t}$. Thus, $\mathcal{F}_{i,t}$ can be expressed in the same form as that of $\mathcal{F}_{i,t}^{S}$ above, removing the term $S_{i,t}$.
We now leverage $\mathcal{F}_{i,t}^{S}$ to conditionally bound $(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}$.
Lemma 12.
Conditional on $\mathcal{F}_{i,t}^{S}$, the term $(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}$ is nonnegative and bounded as
$(\pi_{i,t}-\pi_{i,t}^{\star})x_{i,t}^{\top}\theta_{i,t}^{\star}<\sqrt{\frac{\left\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\right\|^{2}_{\underline{\hbox{\set@color$V$}}_{i,t}}+\beta^{2}_{i,t}(\delta_{1},\delta_{2})\mathbb{E}\left\|\eta_{i,t}\right\|^{2}}{\min(\pi_{\min},1-\pi_{\max})\alpha\min(i,t)}}.$